Session: Feb 2006 (session 1)
Wavelength: 6 cm


To give feedback on this experiment, click

Experiment Feedback

Urumqi (Zhang Hua at Mon Mar 6 06:07:34 2006) Success :

Torun (Kaz Borkowski at Mon Mar 6 10:36:51 2006) Success :
No known problems.

Medicina (C. Bortolotti at Mon Mar 6 11:17:08 2006) Success :
Bad time and wind

Effelsberg (A. Polatidis at Mon Mar 6 11:20:56 2006) Success :

Westerbork (Tony Foley at Mon Mar 6 13:12:10 2006) Minor failures : PHAS
Wb used in phased array mode. Lost from 08:33 to 08:48:30UT with tied-array problems Also note, the last disk pack reported StreamStor1 when it was in fact WSRT-058. I relabelled the pack at the end of the observation. Also some snow (but melted immediately)

Jodrell_Bank (Alastair Gunn at Mon Mar 6 13:13:59 2006) Success :
The MkII telescope was used (Jb2). Note that disk pack OSOD-0058 appears to have failed disks (3/4?) but data recording was fine with reduced capacity. No other known problems.

Cambridge (Alastair Gunn at Mon Mar 6 13:14:54 2006) Success :
Note that Cambridge signal appears in BBCs 3 and 4, and Darnhall appears in BBCs 1 and 2. No known problems.

Hartebeesthoek (Jonathan Quick at Mon Mar 6 13:54:51 2006) Success :
No known problems

Noto (P. Cassaro at Tue Mar 7 16:24:39 2006) Minor failures : WIND
From 08:30 UT observation suspended due to high speed wind.

EVN experiment feedback facility: entire session

Entire Session

Session: Feb 2006 (session 1)
Wavelengths: 6 3.6 5 18 cm


To give feedback on this entire session, click

Entire Session Feedback

Torun (Kaz Borkowski at Mon Feb 20 11:25:29 2006) Minor failures : PCAL
Tr phase cal system failed to produce signals during all the 6 cm experiments for yet unknown reason.

Westerbork (Tony Foley at Fri Feb 24 15:48:24 2006) Minor failures : ANT POIN
The feedback from the antenna control for SINGLE-DISH observations is unreliable for 20seconds after a source change. We will try to fix this. (Affected only the 5cm experiments -Z.P.)